During the General Enhanced Community Quarantine, we were mandated to stay and work from home. We do a lot of walking as we do a lot of household chores at home. Applying vector addition, identify five displacements inside your house then calculate your total displacement. Use component method in finding your resultant displacement. Show your complete solution.


1
Expert's answer
2021-09-20T16:38:42-0400


By vector addition, the displacement is 10 m.

Analytical method:


"D_y=10-10+10=10\text{ m}.\\
D_x=4-4=0.\\
D=\sqrt{D_x^2+D_y^2}=10\text{ m}."